YoursTruly:

Rodrigo is a salesman who earned a base pay of $38,800 and was paid commission of 14% for each book set he sold. If x represents the number of books sold, and y represents his total pay in dollars, then write an equation that best represents Rodrigo's total pay in dollars?

Vocaloid:

Total pay = base pay + (commission/100)*number of books Then replace the words with their matching variable

FOD:

ಠ_ಠ hm.

YoursTruly:

Use the following function rule to find f(4). \[f(x)=(2)^x+6\] \[f(4)=?\]

Zowie:

Okay. This is simple. All you do is plug in. You see in the first equation where the x is? Well you take the 4 and put it where the x is and then solve. Tell me what you get.

YoursTruly:

mkmkmk

YoursTruly:

I got 22

Zowie:

Yew are correct! ^.^ Gewd job babie! <3

YoursTruly:

Ouuui. ;o

Zowie:

Anymoar?

YoursTruly:

I aced them all ;o

YoursTruly:

I might have another math assignment to do that is confusing x'D

Zowie:

Okay. Send then my way.

YoursTruly:

Seema has $20 in a savings account that earns 10% interest, compounded annually. To the nearest cent, how much will she have in 3 years? Use the formula B = p(1 + r)t, where B is the balance (final amount), p is the principal (starting amount), r is the interest rate expressed as a decimal, and t is the time in years. $=

Zowie:

Okay. So we know what the letter (variables) mean (represent). So all you do it plug in the money amounts. B= what we are finding. r= 10% but we have to change to a decimal so it would be .1 because when changing a percent to a decimal you move the . 2 spaces to the right. 10.00 -> .10. t=3 since 3 years. p= $20/20 because that's how much she has right now. So your overall equation is B=20(1+ 0.1)3. Solve that equation and tell me what you get.

YoursTruly:

I got 78

Zowie:

You used PEMDAS right?

YoursTruly:

No. >.> I just typed it into my calculator

Zowie:

You have to use PEMDAS. Retry it using PEMDAS. Tell me what you get then.

YoursTruly:

mk Parenthesis is 1.1 so then it would be 20(1.1)3 right?

Zowie:

No. This si your equation.

Zowie:

B=20(1+ 0.1)3

Zowie:

Oh sorry. I didn't understand. Yes. That's the first step.

YoursTruly:

So when we get to the math do I do 20(1.1) first or (1.1)3 or all together?

Zowie:

Id there is 2 of the same step you do the one that comes first in the equation so you multiply 20*1.1 then 3.

YoursTruly:

Okay It's the same as when I multiply all together. I got 66

YoursTruly:

Okay It's the same as when I multiply all together. I got 66

YoursTruly:

Okay It's the same as when I multiply all together. I got 66

Zowie:

http://prntscr.com/gqcu2y It says you typed it 3 times. FIgured I'd share that with you. Also you are correct! ^.^ <3

YoursTruly:

What the Potato/ x'D It's my mouse. When I click it double clicks sometimes and apparently triple too